OpenStudy (anonymous):

Find the product. a^4(3a^2 - 2a + 1)

OpenStudy (anonymous):

use the distributive property of multiplication.

OpenStudy (anonymous):

a(b+c) = ab+ ac a(b+c+d) = ab +ac +ad Do you know how to solve for it now?

OpenStudy (anonymous):

do I multiply the left, and then add the outcome to the right?

OpenStudy (anonymous):

you just have to multiply a with all the terms in the parentheses, and add those

OpenStudy (anonymous):